OpenStudy (anonymous):

a 3.0kg steel block is a rest on a frictionless surface. a 1.0kg lump of clay is propelled horizontally at 6.0 m/s towards the block. Upon collision the clay and steel block stick together and move to the right with a speed of...?

OpenStudy (kainui):

First off, draw out what you imagine this should look like before and after.  Now we can think about whether or not this is an elastic or inelastic collision. Is momentum conserved or not? Since the surface is frictionless, then momentum is conserved. What is momentum? P = mv So that means the momentum of before collision is the same as after collision. Since the collision is perfectly elastic, then we can treat the "after" part of our equation as a single mass with one velocity. (numbers are subscripts to tell mass 1, the block, apart from mass 2, the clay) m1v1+m2v2 = (m1+m2)v (3.0kg)(0m/s)+(1.0kg)(6.0m/s) = (3.0kg+1.0kg)velocityfinal